# Closure for Microscale Fluctuations in Chaotic Mixing Fronts

### v1

@@ Abstract
Scalar mixing fronts are formed under the stirring and spreading action of a random velocity field. The prediction of concentration statistics below the velocity length scale is an open problem. Here, we show that when fluid agitation is caused by a smooth chaotic flow with weak persistence, the evolution of scalar variance can be uniquely related to the stirring and spreading properties of the flow and the diffusivity of the scalar. Using the spectral description of microscale scalar fluctuations in presence of a large-scale forcing by Kraichnan [1], we uncover the link between variance production and dissipation in a dispersing front. This provides a universal closure for the scalar fluctuation microscale whenever there exists a clear separation between macroscale dispersive forcing and microscale dissipation. We validate the closure model on high-resolution direct numerical simulation of advection-diffusion in chaotic flows of variable strength and persistence. These findings open a new avenue to predict both conservative and reactive transport in incompletely mixed environments, ranging from porous media flows to engineered flows in microfluidic devices.
@@ I Introduction
Figure 1: Scalar front mixed and dispersed by a smooth periodic chaotic flow (sine flow). Scalar concentrations (colorscale) obey Eq. (7) with $\kappa=5\cdot 10^{-6}$ and $\bm{u}$ defined by Eq.(LABEL:eq:sineflow), with $U=0.2$ ($\gamma\approx\sigma^{2}_{\gamma}\approx 0.05$), and $t=300$. All the important length scales appearing in the front mixing problem are shown at real scale in the plot (from small to large). For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$  (1), the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$ (Eq. (28)), the microscale is $s_{m}\approx 0.02s_{v}$ (Eq. (19)), the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ (Eq. (36)) and the dispersive scale $s_{d}\approx 1.41s_{v}$ (Eq. (4)).
@@
Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [2], turbulence  [3], or porous media [4]. Away from the source, mixing fronts naturally develop between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig. 1). Such interfaces are observed, for instance, in the fringe of contaminant plumes in the subsurface [5], in density or thermal currents in lakes or coastal areas, or in river junctions [6]. A universal modelling framework that captures the mixing process in such interfaces is still missing, even in the simple case where the solute is passively transported by the flow and the flow is smooth at the scale of interest. In contrast to rough turbulent flows, smooth chaotic flows are characterized by a minimum velocity lengthscale $s_{v}$ below which velocity gradients are constant (Fig. 1). Such conditions are found for laminar flows at low Reynolds numbers or in turbulent flows at high Schmidt numbers, where dissipative scales are larger than mixing scales. Below $s_{v}$, solute mixing occurs by the simultaneous action of fluid deformation and molecular diffusion. The Batchelor scale [7] is the minimum lengthscale at which scalar fluctuations can persist (Fig. 1), given by the length scale at which fluid deformation balances by molecular diffusion as
$s_{B}=\sqrt{\kappa/\gamma},$ (1)
with $\kappa$ the molecular diffusivity and $\gamma$ the mean stretching rate of the chaotic flow. The value of $\gamma$ is dictated by the accumulation of fluid deformations produced by microscale velocity gradients, scaling as $U/s_{V}$ where $U$ is a characteristic velocity scale. When the characteristic persistence time $T$ of velocity gradients is small compared to the characteristic advection time on the velocity scale $s_{V}/U$, the mean stretching rate is expected [1] to scale as
$\gamma\sim T\,(U/s_{v})^{2}.$ (2)
Flow persistence can be quantified by a dimensionless persistence parameter, or Kubo number [8, 9],
$Ku=\frac{UT}{s_{V}}.$ (3)
For $Ku\ll 1$, the flow has weak persistence while for $Ku\gg 1$, the flow has long time correlations. Eqs.(1) and (2) show that scalar fluctuations are maintained below $s_{v}$ whenever ${\kappa}<U^{2}T$, that is, for weakly diffusing or strongly agitated solutes. In unbounded domains, chaotic flows favour the dispersive spreading of solutes leading to the formation of scalar gradients at increasingly large scales
$s_{d}\sim\sqrt{Dt}$ (4)
where $D$ is the dispersion coefficient and $t$ is the time (Fig. 1). Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
$D\sim U^{2}T.$ (5)
Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are constant, we have the general lengthscale separation
$s_{d}\gg s_{v}>s_{B}$ (6)
at least asymptotically in time. Practically, resolving scales smaller than $s_{d}$ or $s_{v}$ is often impossible numerically at small $\kappa$, due to the large number of grid cells necessary to resolve $s_{B}$. Thus, one aims at finding macroscale equations that correctly capture small-scale fluctuations into an effective model. In the majority of chaotic flows, such effective description remains largely unknown.
At the microscale, the lamellar theory [10, 11] accurately predicts the evolution of elementary solute filaments, which elongate and dilute by the coupled action of flow stretching and molecular diffusion. In this one-dimensional representation, the filaments are considered to be independent of each other. Thus, the theory does not allow prediction of mixing statistics in the late time when filaments coalesce and where the concentration field is not Gaussian. Aggregation theories, whether considered above[12] or below [13] the velocity length scale $s_{v}$, have attempted to describe the late-time evolution of the scalar statistics through the evolution of the number of filament coalescence. Well constrained in confined systems [14], aggregation theories are difficult to apply in dispersing interfaces because the density of lamella aggregation becomes a function of both space and time  [15]. For chaotic flows in particular, the exponential increase of material lengths in slowly increasing dispersive volumes (Eq.(4)) questions the applicability of a uniform aggregation process [16], since it would lead to the removal of all scalar fluctuations inside the dispersion zone at a finite time. In contrast, we expect chaotic agitation to maintain fluctuations in proportion to the mean concentration gradients (Fig. 1). In particular, numerical simulations Tsang et al. [17], Haynes and Vanneste [18] in periodic domains have shown that the decay rate of scalar variance is controlled by microscale dissipation processes in domains smaller than a few $s_{v}$, while it is controlled by large-scale gradients on larger domains.
Batchelor [7], Kraichnan [1] and later Kalda [19] proposed an alternative description of mixing by smooth chaotic flow, recognising that concentration gradients are transported to smaller scales by random fluid stirring, while molecular diffusion is only effective at small scales to reduce scalar variance. These authors describe this process through an evolution equation for the power density spectrum (PSD) of scalar fluctuations as a function of time and spatial frequency. While relying on a similar description of the fluid deformation process as the lamellar representation, the spectral description is global in that it is not limited to isolated Gaussian lamellae. Thus, it is well adapted to the case of scalar mixing fronts of arbitrary shape, where single lamellar elements are not distinguishable. To the author’s knowledge, a comparison of such a spectral description with direct numerical simulations of scalar mixing in smooth chaotic flows has never been attempted.
While appealing, the spectral description of mixing does not directly inform on the concentration statistics of the mixture, an information which is critical to many transport applications, e.g. for contaminant spreading or reactive transport. Furthermore, there is currently no consensus on the asymptotic shape of the scalar Probability Density Function (pdf) in smooth chaotic flows, with experimental studies pointing to Gamma [20], Exponential [4], Power-law [17, 21] or even Gaussian pdfs [22]. Theoretical results on the advection-diffusion equation [23] constrain the possible shapes of the scalar pdf, but leave open the underlying physical mechanisms. In particular, the relation between the asymptotic scalar pdf moments of a mixing front and the characteristics of the microscale stirring process has not been not established. In this study, we show that scalar fluctuations within a mixing front are controlled both by the stretching properties of the microscale velocity field and by the dispersive spreading of the scalar. Using Kraichnan [1] model, we propose a physical-based closure for microscale fluctuation and validate it against direct numerical simulations of chaotically agitated fronts.
@@
The paper is organised as follows. In section 2, we recall the theoretical foundations allowing to predict the evolution of scalar variance in mixing fronts and the spectrum of microscale scalar fluctuations. Based on these two elements, we propose a theoretical closure for the scalar dissipation. In section 3, we compare these analytical predictions with direct numerical simulations. We consider a smooth and single scale flow field, the random sine flow, as a prototype of chaotic flow for which the persistence, the dispersivity and the stretching rate can be varied. Sine flows have been extensively used to uncover the key feature of chaotic mixing [24, 25]. Their relative simplicity makes them computationally tractable even for large domains and times. We numerically study the evolution of the concentration fluctuation spectrum in the case of the forcing of large-scale scalar gradient. The numerical spectrum shows the characteristic $k^{-1}$ scaling between $s_{d}$ and $s_{B}$, and accurately matches the Kraichnan’s prediction. We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately capture the evolution of mixing fronts. We finally conclude on the potential of our results to model other transport processes.
@@ II Theory
In this section, we recall earlier results on the behaviour of concentration fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations. Based on these two elements, we then propose a closure model for microscale fluctuations in the presence of a mean scalar gradient.
@@ II.1 Background
II.1.1 Concentration microscale
@@
where $\bar{\bullet}$ is the ensemble mean of the microscale chaotic agitation process, of length-scale $\leq s_{v}$. $\bar{\bullet}$ can also be interpreted as the result of a low-pass filtering with characteristic scale $s_{v}$.
Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$, such that the fluctuation satisfies the inhomogeneous ADE
where the source $f$ has zero mean. For temporally-periodic flows with flow period $T$, the fluctuation $c^{\prime}$ in a zero mean field ($\bar{c}=f=0$) is comprised of a finite set of $N$ strange eigemmodes [26, 21, 27], which are Floquet modes of the AD operator $\mathcal{L}$ of the ADE (7) that manifest as exponentially decaying patterns that control the mixing process:
$c^{\prime}(\mathbf{x},t)=\sum_{k=0}^{K}\alpha_{k}(t)\varphi_{k}(\mathbf{x},t)e^{-\lambda_{k}t},$ (11)
where the (possibly complex) eigenmode $\varphi_{k}(\mathbf{x},t)$ is $T$-periodic in time, and the eigenvalue $\lambda_{k}$ (with $|\lambda_{k}|\leqslant|\lambda_{k+1}|$) controls the persistence of the mode. The coefficient $\alpha_{k}$ is governed by the initial condition $c^{\prime}(\mathbf{x},0)$ as is time-dependent due to the non-self adjoint nature of $\mathcal{L}$, where spectral energy is transferred toward the lowest energy (low $k$) modes with time [28]. With time, only the dominant non-trivial mode $\varphi_{0}$ persists, leading to the observation [21, 26] of convergence of $c^{\prime}$ to a persistent yet decaying spatio-temporal pattern in the case of time-period flows. Note strange eigenmode analogues also exist for aperiodic and open flows [29]. In the presence of a source term such as that given by the mean field in (10), the initial conditions for $c^{\prime}$ decay and the fluctuation converges [30] to
[…]
where $p_{k}(t)$ is also $T$-periodic and typically $|p_{k}|\geq|p_{k+1}|$, hence the low-$k$ modes dominate the fluctuation. Hence, the fluctuation $c^{\prime}$ admits truly persistent spatio-temporal patterns due to continual excitation of the natural mixing modes $\varphi_{k}$ by the mean field $\bar{c}$. Equation (12) shows how the transfer of scalar energy between the mean field and small-scale fluctuations shapes the asymptotic structure of the fluctuation and how the natural mixing modes act to dissipate this scalar energy. Although Eqs. (11)–(12) refer to the concentration field in the real space, we will see below that they have direct analogues in the spectral space. The fluctuations $c^{\prime}$ have zero mean and variance
$\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}},$ (13)
and the mean concentration obeys the isotropic advection dispersion equation at macroscale [31],
@@
When microscale fluctuations are isotropic, and so may be expressed as $\bm{D}=D\mathbf{I}$ The dispersive limit is reached for any velocity fluctuations that have a finite correlation time and integrable statistics [31, 33]. Multiplying Eq. 7 by $c$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations [33, 32]
@@
The left hand side is the transport of scalar variance through advective and dispersive fluxes while the right-hand side is the dissipation and production of scalar variance due to molecular diffusion and dispersive motion respectively. As shown by Eq.(16), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other. In contrast, the destruction of the scalar variance results from microscale diffusive fluxes. Dimensional arguments imply that dissipation can be expressed as
$\overline{(\nabla c^{\prime})^{2}}\sim\sigma^{2}_{c}/s_{m}^{2},$ (17)
where $s_{m}$ is a concentration microscale [34] characteristic of the microscale mixing process. Providing an expression for $s_{m}$ is the main closure problem. Two limiting scenarios can be distinguished depending on the presence of a large-scale forcing process. The first is the decaying scenario where the scalar homogeneizes according to (7), in the absence of any source or boundary conditions. In that case, there is no production of scalar variance and assuming negligible transport,
$\sigma^{2}_{c}\sim\exp{(-2\kappa\,t/s_{m}^{2})},$ (18)
i.e., an exponential decay of scalar variance.The second case arises in the presence of a mean scalar gradient, for instance in a mixing front which is slowly evolving compared to the microscale agitation process. In this case, the left-hand side of Eq. (16) can be neglected, and there is equilibrium between dissipation and production terms. The scalar variance reaches a steady-state with
$\sigma^{2}_{c}=\frac{D}{\kappa}s_{m}^{2}(\nabla\bar{c})^{2}.$ (19)
Large scale concentration gradients spontaneously develop in unbounded mixing domains, due to dispersive effects (Eq. (42)). Thus, the second scenario is the most general and so is considered herein. Since the minimum scale of fluctuations in chaotic flows is fixed on the Batchelor scale, we may first argue that
$s_{m}\sim s_{B}.$ (20)
Inserting this closure in Eq. (19) leads to $\sigma^{2}_{c}=D(\nabla\bar{c})^{2}/\gamma$, independent of $\kappa$. In contrast, numerical simulations [18] generally show a weak dependence on $\kappa$, suggesting that a closure based on the Batchelor scale only is not satisfactory. In the following, we show instead that the transfer of energy from large to small scales globally controls the amount of scalar variance residing in mixing fronts.
II.1.2 Microscale power spectra
As noted by Batchelor [7], the stretching action of the velocity field transfers scalar fluctuations from larger length-scales toward $s_{B}$. The energy spectrum of concentration fluctuations is $E_{k}(k)=|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
@@ II.1.2 Microscale power spectra
with $k=|\bm{k}|=2\pi/\lambda$ the wavenumber, inversely proportional to the wavelength $\lambda$. The scalar spectrum was shown to present a distinct shape in the two mixing scenarios considered above. In the case of unforced scalar decay, the spectrum is essentially flat above the dissipation scale [17]. In the case of forced mixing (for example, in the presence of a large-scale concentration gradient), the energy spectrum shows a characteristic $E_{k}\sim k^{-1}$ scaling below the injection scale (for $k>2\pi/s_{v}$), suggesting the transfer of large-scale fluctuations to small scales by the stirring action of the flow [7]. Above the injection scale, the spectrum must be essentially Gaussian, reflecting dispersive processes (Eq. (14)). Intermediate cases with $k$ exponents between 0 and -1 were found in numerical simulations with periodic domains with periodicity comparable with $s_{v}$[17, 35] suggesting the transition between micro- and macro-scale mixing mechanisms. The evolution of the scalar spectrum of $c^{\prime}$ was related to the flow stretching statistics by Kraichnan [1] and independently by Kalda [19] in the case of a $\delta$-time correlated random velocity field. The theory rests on the fact that, in the absence of diffusion, the scalar gradients follow the same evolution as the elongation of material elements $\rho=\ell/\ell_{0}$ (or material surfaces in 3D), with $\ell$ and $\ell_{0}$ the final and initial length (or area in 3D) of a fluid element, respectively, which increases due to fluid deformation. The theory is based on three main elements. The first element is that in a $\delta$-correlated flow, the evolution equation for the pdf of elongation $\Lambda(\ln\rho)$ is
$\frac{\partial\Lambda}{\partial t}+\gamma\frac{\partial\Lambda}{\partial\ln\rho}=\sigma^{2}_{\gamma}/2\frac{\partial^{2}\Lambda}{\partial\ln^{2}\rho}.$ (22)
where $\gamma$ is the mean stretching rate and $\sigma^{2}_{\gamma}$ the variance of the stretching rate [19, 36] defined as
$\gamma$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{\ln\rho(t)},$ (23)
$\sigma^{2}_{\gamma}$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{(\ln\rho(t)-\gamma)^{2}}.$ (24)
Eq. (22) shows that $\Lambda$ evolves according to a mean drift with the mean stretching rate while broadening due to fluctuation in stretching. Note that in $\delta-$correlated chaotic flows, the stretching variability is related to the dimension of the space $d=2,3,\cdots$ by  [1, 25]
@@
The second element is that an increase in $\ln\rho$ induces the same increment of wavenumber $\ln k$ in the spectral concentration gradient domain, because fluid stretching is compensated for by transverse compression, which transfers scalar gradients to smaller wavelengths. Thus, in absence of molecular diffusion, the power spectra density of scalar gradients $\Phi(k)\sim E_{k}k$ must follow Eq. (22), with $\ln\rho$ replaced by $\ln k$.
The last element is that molecular diffusion becomes efficient at large $k$ to dissipate energy at a rate $2\kappa k^{2}$, which can be included in the evolution equation as
$\frac{\partial\Phi}{\partial t}+\gamma\frac{\partial\Phi}{\partial\ln k}=\frac{\sigma^{2}_{\gamma}}{2}\frac{\partial^{2}\Phi}{\partial\ln^{2}k}-2\kappa k^{2}\Phi.$ (26)
In essence, such model resembles the lamellar description of mixing [37] with the difference that diffusion now operates on an arbitrary concentration field rather than single Gaussian lamellae. This equation has been studied by Kraichnan [1] and Kalda [19]. In the case of a source of large-scale concentration gradients, the boundary condition is $\Phi(0,t)\equiv\chi_{0}/\gamma$, where $\chi_{0}$ is the rate at which scalar variance is injected into the system from these large-scale gradients and $\gamma$ is the mean stretching rate of an infinitesimal line element [7]. Kalda [19] provides an explicit solution to Eq. (26) in terms of Hankel functions. Kraichnan [1] proposes a simpler equivalent form (5.14 in the original paper)
$E_{k}(k)=\frac{\chi_{0}}{\gamma k}f(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (27)
where the length scale
$s_{B}^{\prime}=\sqrt{\kappa/\sigma^{2}_{\gamma}}.$ (28)
is a modified Batchelor scale, and $f(k)$ is the solution of the Kummer’s equation [1]
$kf^{\prime\prime}-(2k+2\gamma/\sigma^{2}_{\gamma}-1)f^{\prime}+(2\gamma/\sigma^{2}_{\gamma}-1)f=0,\quad f(0)=1.$ (29)
As shown by Kraichnan [1], in two dimensions, $\gamma/\sigma^{2}_{\gamma}=1$, and $f\propto k^{1/2}$ at large $k$.
Equation (27) describes scalar fluctuations in the Batchelor regime, that is, when $\gamma$ is not scale-dependant, and shows that for low wavenumbers the energy decays as $k^{-1}$, with an exponential cutoff at high wavenumbers. In contrast, above the characteristic velocity length scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq. (16)). In the following, we show that the two descriptions (chaotic and dispersive) can be matched on a particular injection scale $s_{i}$, scale at which the scalar variance is transmitted at higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
II.2 Proposed microscale closure
As shown in Eq.(16), scalar fluctuations are produced at large scale by the mean dispersive flux. We define the injection scale $s_{i}$ as the characteristic scale at which scalar fluctuations are produced inside the mixing front. In turn, once produced by dispersive motions, scalar fluctuations are transferred to smaller scales by the cascade process described by Kraichnan [1](Eq. (27)). The amount of scalar variance residing in the system between the injection and the dissipation scale (Eq. (13)) can be obtained by integration, e.g.
$\sigma^{2}_{c}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (30)
where $E_{k}$ is given by Eq. (27). The integral is analytical when $f=1$, and gives
where $Ei$ is the exponential integral. We show in Fig. 6 that Eq. (31) accurately describes numerical simulations, even in the case of two-dimensional flows where $f\neq 1$. This is explained by the primary importance of the exponential cutoff in setting the value of the integral (34), rather than the exact shape at intermediate wavenumbers.
The value of $\chi_{0}$ is controlled by the large-scale dispersion process. At large scale, the variance production rate due to mean scalar gradients (Eq. (16)) is:

### v2

# Mixing Fronts in Chaotic Flows

@@ Abstract
Mixing fronts develop at the interface of fluids with different solute concentrations leading to jointly evolving macroscopic and microscopic concentration gradients. Macroscopic gradients decay along the front through dispersion driven by microscopic velocity heterogeneity. The resulting macroscopic fluctuations are transmitted at the micro-scale through stretching and compression, where they are ultimately dissipated. While the elementary mechanisms leading to dispersion on the one hand and scalar mixing on the other hand are well understood, predicting how their coupling governs the evolution of concentration statistics within dispersing fronts remains a challenge. Here, we use theoretical derivations and numerical simulations of mixing fronts in chaotic flows to link the evolution of scalar variance to the microscale stirring and macroscale spreading properties of the flow. We argue that the transfer of energy between the macroscopic and microscopic scales operates at a characteristic length scale $s_{i}$, that we derive by comparing the characteristic times for scalar variance decay from dispersion and mixing. This leads to a closed expression for the concentration variance, which captures the results of numerical simulations with no fitting parameters for a broad range of Peclet numbers. For increasing stirring persistence, the numerical simulations deviate from the theory in the small Peclet range, which we explain qualitatively. These findings open a new avenue for predicting both conservative and reactive transport in mixing fronts, ranging from porous media flows to engineered flows in microfluidic devices.
@@ I Introduction
Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [1], turbulence  [2], or porous media [3]. Away from the solute source, mixing fronts naturally develop at the interface between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig. 1). Such interfaces are observed, for instance, in the fringe of contaminant plumes in laminar flows through porous media [4], or in turbulent density or thermal currents in lakes, coastal areas, or river junctions [5]. Describing concentration fluctuations in mixing fronts is of particular importance in the context of mixing-driven reactions in laminar [6] or turbulent chaotic flows [7]. Turbulent and laminar chaotic flows differ by the presence or absence of wide spectrum of velocity gradient scales.
Figure 1: Scalar front mixed and dispersed by a single-scale periodic chaotic flow (sine flow). Scalar concentrations (colorscale) obey Eq. (1) with $\kappa=5\cdot 10^{-6}$ and $\bm{u}$ defined by Eq.(LABEL:eq:sineflow), with $U=0.2$ ($\gamma\approx\sigma^{2}_{\gamma}\approx 0.05$), and $t=300$. The mixing scales appearing in mixing fronts are shown at real scale in the plot (from small to large). For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$, the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$, the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ and the dispersive scale $s_{d}\approx 1.41s_{v}$.
While classical scenarios of forced turbulent mixing in closed domains [8, 9, 10], or solute dilution close to a solute source [11] have been well studied, mixing in dispersing fronts [12] is less understood. In scalar turbulence, most existing models have focused on “homogeneously heterogeneous” mixing conditions [8], or micromixing, where the existence of macroscale dispersive transport, or macromixing, is neglected [7]. In turbulent scalar plumes, Baldyga and Bourne [12] suggest that there exists a competition between macro and micro-mixing phenomena in the presence of large scale gradients. In laminar chaotic flows with a single velocity length scale $s_{v}$ (Fig. 1), scalar mixing is well predicted below $s_{v}$ by the local mechanism of stretching-enhanced diffusion [13]. At the macroscale (above $s_{v}$), it is also well-known that dispersive transport globally controls scalar variance in bounded domains [14, 15]. However, how these two processes, operating at different scales, dictate the evolution of concentration statistics in dispersing scalar fronts is not fully understood  [12].
Although mean concentration fields are generally well described by effective Fickian dispersion models [16], a description of scalar fluctuations generally requires additional closure arguments. Micromixing closure arguments can be sought in the lamellar theory [17, 18], which predicts the enhancement of scalar gradients by the stretching of isolated solute filaments and, notably, the mixing time after which scalar dissipation is effective. Aggregation theories [19, 20] can be used to describe the evolution of concentration statistics in the regime where solute filaments overlap with each other [21]. However, the aggregation formalism is difficult to apply to dispersing mixing fronts because aggregation becomes a function of both space and time [22]. Mixing theories describing the evolution of the scalar power-spectra in response to smooth fluid deformations have shown good accuracy in capturing microscale scalar mixing in bounded domains, in both the case of strong [23] or weak [24] stretching persistence, within the use of only a few numerical constants [10]. However, they have not yet been applied to the case of freely dispersing fronts.
In this study, we investigate the relation between macroscale dispersive transport processes above $s_{v}$ and microscale mixing processes below $s_{v}$ in laminar single-scale chaotic flows. We use the spectral description of microscale mixing to provide a closure for scalar fluctuations within a dispersing front mixed by a smooth chaotic flow. We argue that there exists a fixed injection scale $s_{i}$, at which macroscale scalar fluctuations controlled by dispersive spreading are transferred to the microscale, and dissipated by stretching-enhanced diffusion. This scale is set by the balance between macroscale dispersive spreading and microscale chaotic mixing. This leads to a prediction of the concentration variance within mixing fronts with no fitting parameters.
@@
The paper is organised as follows. In section 2, we recall the equations governing the conservation of scalar variance in mixing fronts and the spectral evolution of microscale scalar fluctuations. Based on these two elements, we propose a theoretical closure for the scalar dissipation. In section 3, we compare these analytical predictions with direct numerical simulations. We consider a smooth and single scale flow field, the random sine flow, as a prototype of chaotic flow for which the persistence, the dispersivity and the stretching rate can be varied. Sine flows have been extensively used to uncover the key feature of chaotic mixing [25, 26]. Their relative simplicity makes them computationally tractable even for large domains and times. We numerically study the evolution of the concentration fluctuation spectrum in the case of the existence of a large-scale scalar gradient. We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately captures the evolution of mixing fronts. We finally conclude on the potential applicability of our results to model other transport processes.
@@ II.1 Background
In this section, we recall earlier results on the behaviour of scalar concentration mean and fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations. Based on these two elements, we then propose a closure model for microscale fluctuations in the presence of a mean scalar gradient.
II.1.1 Macroscale evolution of concentration mean and variance in a dispersing front
@@
where $\bar{\bullet}$ is the ensemble mean of the microscale chaotic agitation process, of length-scale $\leq s_{v}$. $\bar{\bullet}$ can also be interpreted as the result of a low-pass filtering with characteristic scale $s_{v}$. Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$. Inserting (2)-(3) in (1), and averaging gives
@@
When microscale velocity fluctuations are isotropic, the dispersion tensor is also isotropic and expressed as $\bm{D}=D\mathbf{I}$. Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
$D\sim U^{2}T.$ (6)
with $U$ and $T$ a characteristic flow agitation velocity and time. The dispersive limit is reached for any chaotic flow of finite correlation time and integrable statistics [16, 28]. Dispersion leads to the spreading of solute and the formation of scalar gradients at increasingly large scales $s_{d}\sim\sqrt{Dt}$ in mixing fronts (Fig. 1).
The fluctuating part of the concentration field satisfy the inhomogeneous ADE
where the source $f$ has zero mean. Multiplying Eq. 7 by $c^{\prime}$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations $\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}}$  [28, 27]
@@
The left hand side is the transport of scalar variance through advective and dispersive fluxes while the right-hand side is the dissipation and production of scalar variance due to molecular diffusion and dispersive motion respectively. As shown by the right-hand side of Eq.(8), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other. In contrast, the destruction of the scalar variance results from microscale diffusive fluxes.
The objective of the paper is to provide a physically-based closure for microscale scalar dissipation in the case of a freely dispersing front mixed by a smooth chaotic flow.
II.1.2 Microscale power spectra in a smooth chaotic flow
As noted by Batchelor [23], the stretching action of a smooth chaotic flow transfers scalar fluctuations from $s_{v}$ towards the so-called Batchelor scale $s_{B}=\sqrt{\kappa/\gamma}$, where $\gamma$ the mean stretching rate of the chaotic flow. The value of $\gamma$ is dictated by the accumulation of random fluid deformations produced by microscale velocity gradients, scaling as $U/s_{v}$ where $U$ is a characteristic velocity scale.
Kraichnan [24] described the evolution of the scalar power spectrum (PSD) in the case of weak chaotic flow and in the absence of macroscopic concentration gradients. Flow persistence can be quantified by the Kubo number [29, 30], $Ku={UT}/{s_{v}}$, where $T$ is the characteristic persistence time $T$ of velocity gradients. When $Ku\ll 1$, the mean stretching rate is expected to scale as
$\gamma\sim T\,(U/s_{v})^{2}.$ (9)
In addition, analytical results can be obtained on the shape of the microscale PSD $E_{k}(k)\equiv|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
@@ II.1.2 Microscale power spectra in a smooth chaotic flow
with $k\equiv|\bm{k}|=2\pi/s$ the wavenumber, inversely proportional to the wavelength, or scale $s$. In the case of a large scale source of scalar fluctuations, Kraichnan [24] showed that the PSD follows
$E_{k}(k)=\frac{\chi_{0}}{\gamma k}{F}_{\gamma/\sigma^{2}_{\gamma}}(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (11)
where $s_{B}^{\prime}\equiv\sqrt{\kappa/\sigma^{2}_{\gamma}}$ is a modified Batchelor scale defined with the variance of the stretching rate $\sigma^{2}_{\gamma}$ rather than the mean $\gamma$, and the function ${F}_{\gamma/\sigma^{2}_{\gamma}}(k)$ is the solution of a second-order ODE with parameters $\gamma$ and $\sigma^{2}_{\gamma}$ (see Supp. Mat. for details and Eq. (5.14) in Kraichnan [24]). In weakly persistent flows, $\sigma^{2}_{\gamma}$ is related to $\gamma$ through the dimension of the space $d$ where mixing occurs [24, 26]:
@@
In particular, when $d=3$, $\gamma=3\sigma^{2}_{\gamma}/2$ and ${F}_{3/2}=1$. In turn, when $d=2$, $\gamma=\sigma^{2}_{\gamma}$, $s_{B}=s_{B}^{\prime}$, and ${F}_{1}\approx 1+O(k^{1/2})$ [24]. Note that both $\sigma^{2}_{\gamma}$ and $\gamma$ have units of time frequencies. Eq. (11) shows that the PSD decays as $k^{-1}$ for low wavenumbers, with an exponential cutoff at high wavenumbers. Note that the Kraichnan spectrum is comparable to Batchelor’s prediction [23] obtained in the case of a quasi-static (long persistence) straining motion, with the difference that the latter has a steeper cutoff ($E_{k}\sim\exp(-k^{2}s_{B}^{2})$) at high wavenumbers due to the absence of stretching fluctuations.
The Kraichnan [24] theory assumes that the stretching field is smooth ($\gamma,\sigma^{2}_{\gamma}$ are constants independent of scale), thus only capturing scalar fluctuations at scales smaller than $s_{v}$. Hence, it does not describe how these microscale fluctuations interact with macroscopic concentration gradients at large scale ($s_{d}\gg s_{v}$). A this scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq. (8)). The low wavenumber PSD of freely dispersing fronts is thus expected to follow
$E_{k}(k)\sim\exp(-2Dk^{2}t),$ (13)
. Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are fixed, dispersing fronts show the characteristic separation of lengthscales
$s_{d}\gg s_{v}\gg s_{B},$ (14)
at least asymptotically in time. In the following, we investigate how to link the Kraichnan description of small scale concentration fluctuations with the macroscopic evolution of large scale concentration gradients in mixing fronts. We show that the descriptions of mixing at macroscale (Eq. (8)) and at microscale (Eq. (11)) can be matched on a particular injection scale $s_{i}$, at which the macroscale scalar fluctuations are transmitted to higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
II.2 Scalar variance in mixing fronts
Since scalar fluctuations in a freely dispersing front are produced at large scales by the mean dispersive flux (Eq.(8)), we assume that the rate of variance injection $\chi_{0}$ is controlled by the large-scale dispersion process. This hypothesis is similar to the spectral equilibrium hypothesis in turbulence, which assumes that the scalar dissipation rate is set by the inertial-convective range, regardless of the extent of the viscous-diffusive range [7]. The variance production rate due to mean scalar gradients is: